OpenStudy (drshuler23):

If , what is the average rate of change of f(x) over the interval [1, 5]?

OpenStudy (hedgepig):

We're missing some information here.

OpenStudy (drshuler23):

f(x)=x^3-3

OpenStudy (hedgepig):

And what's the formula for average rate of change?

OpenStudy (drshuler23):

A(x)=f(b)-f(a)/b-a

OpenStudy (hedgepig):

and what numbers are we given?

OpenStudy (drshuler23):

A) 29.5 B)30 C)43 D)120

OpenStudy (hedgepig):

No, I mean, in the original question, we are given 1 and 5. Where do they go in the formula?

OpenStudy (drshuler23):

this is the original question: If f(x)=x^3-3, what is the average rate of change of f(x) over the interval [1, 5]?

OpenStudy (hedgepig):

Yeah - so what do we do with the 1 and the 5 inside the formula?

OpenStudy (drshuler23):

i dont know